# Many-Body Dissipative Particle Dynamics Simulations of Lipid Bilayers with the MDPD-MARTINI Force-Field

### v1

@@ Abstract
Many-body dissipative particle dynamics (MDPD) offers a significant speed-up in the simulation of various systems, including soft matter, in comparison with molecular dynamics (MD) simulations based on Lennard-Jones interactions, which is crucial for describing phenomena characterized by large time and length scales. Moreover, it has recently been shown that the MARTINI force-field coarse-graining approach is applicable in MDPD, thus rendering feasible the simulation of complex systems as in MD MARTINI for ever larger systems for longer physical times. Here, simulations of various lipid membranes were performed by using the MDPD-MARTINI coarse-grained (CG) force-field, relevant properties were calculated, and comparison with standard MD MARTINI CG simulations and experimental data was made. Thus insights into structural properties of these bilayer systems and further evidence regarding the transferability of the MDPD-MARTINI models are provided. In this regard, this is a natural next step in the development of the general-purpose MDPD-MARTINI CG force-field, which generally provides significant speed-ups in both computational and physical simulated times, in comparison with standard CG MD simulations.
@@
Many-Body Dissipative Particle Dynamics, Molecular Dynamics, MARTINI Force-Field, Lipid Bilayers
preprint: Preprint
I Introduction
@@
Molecular dynamics (MD) [1] is a common method used for the simulation of a whole range of materials in various areas, such as soft matter, fluid dynamics, and solid-state physics. A key element for reliable MD simulations is the ability of describing the interactions between particles (force field) in such a way that various experimental properties of interest of a system are well-reproduced by the simulation. To this end, the non-bonded interactions between particles play a key role in reproducing these properties and much of the research for creating the various force-fields (FFs) is mainly focused on adequately describing those non-bonded interactions for a given cutoff at a specific temperature (the development of FFs valid over a wide range of temperatures is an active field of research), with further refinements coming from the topology FF parameters. When interactions are considered between all atoms of the system, that is in the case of all-atom FFs, simulations are generally computationally expensive, albeit often necessary, for example, when a finer time- and length-scale resolution is required, what has led to the development of coarse-grained (CG) FFs [2, 3, 4, 5]. In this case, interactions between group of atoms are modeled instead of individual atoms, thus the number of required force (interactions) calculations during the integration of the equations of motion in MD simulations is significantly reduced, which, in turn, results in faster simulations and significant savings in energy resources consumed to run the simulations and cool down the hardware. Moreover, CG models offer the possibility of using a larger time step in the simulation and faster dynamics, which provides an additional speed-up of the simulations.
@@
MARTINI is one of the most popular CG FFs for MD simulation, since it is a general-purpose FF that can be applied in the simulation of a wide spectrum of components [6, 7, 8, 9], such as proteins [10, 11, 12, 13], polymers [14], carbohydrates [15], glycolipids [16], glycans [17], DNA [18], RNA [19], water [20] and various solvents [21]. In addition, various extensions include simulations for specific pH [22] and more recently chemical reactions (reactive MARTINI) [23]. While the MARTINI 3.0 FF [24] has recently been released, the development of any FF is a continuous effort that aims at improving the FF parameters through refinement based on a wider range of methodologies and validation with experimental data on an ever growing number of possible systems. In the case of CG FFs, validation can often take place by comparison to results obtained by well-established all-atom FFs or CG models.
@@
Regarding the MARTIN MD FF, its popularity is mainly due to the fact that it allows for the simulation of a wide spectrum of different systems by only using a finite number of CG bead types and a top-down approach for non-bonded interactions. Here, the concept of a ’LEGO’ approach of combining these beads to build more complex molecules while accounting for chemical specificity is used. In addition, MARTINI is based on the Lennard-Jones potential with the smallest possible cutoff including attractions, which, in principle, renders this model computationally efficient, while various features implemented in GROMACS software (e.g. virtual atoms) [25], namely the native platform for MARTINI, allow for further speed-up of the simulations. Currently, the MD MARTINI method enjoys a large community, which applies and further tests the models in a range of diverse systems, thus practically participating in the validation effort of this general-purpose FF.
@@
Despite the wide use of the MD MARTINI model, MD is still a computationally expensive method, even for CG models. However, it has recently been shown that the MARTINI approach can be applied in the case of many-body dissipative particle dynamics (MDPD) [26], a method that offers a computational speed-up of about an order of magnitude in comparison with the MD method due to the short-range interactions, as well as an additional speed-up from the use of soft interactions, while at the same time it accounts for both repulsion and attraction between particles [27, 28, 29, 30, 31, 32]. Moreover, there is no limitation in combining MDPD and MD models in the same simulation. However, such a scenario is in general undesirable, since it will naturally incur a higher computational cost for the simulation. Here, we build on previous work that dealt with the study of DPPC bilayers and further apply the MDPD-MARTINI FF in investigating POPC and DOPC lipid bilayers and compare the obtained results with those of standard MD-MARTINI simulations. Such a comparison between the MD and MDPD MARTINI FFs is realized on the basis of a number of related properties for bilayer systems. Moreover, a new interaction level is provided to account for the C3 MARTINI bead-type in DOPC and POPC lipids. The MDPD-MARTINI models provides equivalent results to the MD MARTINI models, while they offer at the same time a significant speed-up in the simulations, namely several minutes of MDPD simulations instead of several hours of MD simulations, while interactions are transferable as in the case of MD MARTINI.
@@
II Simulation models and methods
Figure 1: MD and MDPD models of DPPC, DOPC and POPC lipids. The type of beads are indicated by the different color. In particular, $Q_{0}$ represents the choline headgroup, $Q_{a}$ the phosphate headgroup, two $N_{a}$ beads correspond to the glycerol ester moiety, while the two alkane chains consist of $C_{1}$ and $C_{3}$ (for POPC and DOPC) beads. The POPC has a double bond at carbon atom positions 9 and 10, while the DOPC has a double bond at each chain, as illustrated. The interactions between the beads are given in Table 2.
II.1 Molecular Dynamics
The MD model follows that of previous work for DPPC bilayers [26], which has been a benchmark system in the case of the MD MARTINI FF [33] and therefore further details will be here omitted. Initial files for LAMMPS software [34, 35] for the DPPC, DOPC, and POPC bilayers based on the MD-MARTINI model can be found as moltemplate examples [36]. The DPPC model consists of different bead types, that is $Q_{0}$ for the choline headgroup, $Q_{a}$ for the phosphate headgroup, two $N_{a}$ beads for the glycerol ester moiety, and two chains, each comprised of four $C_{1}$ beads, to represent the alkane tails. For DOPC and POPC lipids, the model differs from that of DPPC lipids. This distinction arises due to the presence of double bonds in the alkane chains, with one double bond in POPC and two in DOPC. These double bonds require a different angle in the chain structure and an additional bead type, C3, to accurately represent their molecular geometry The lipid models are illustrated in Figure 1, with the names of the bead types being the same for both the MD and MDPD MARTINI models following the MARTINI FF naming convention. Hence, this implies that both models use the same mapping of CG beads to chemical group of atoms. Finally, water is represented by the P4 bead type.
@@ II.1 Molecular Dynamics
Typical simulations for forming the lipid bilayers follow the same protocol as in previous work [26]. However, here, the aim is not to show the computational speed-up of the method during the bilayer formation, since this is well-established for MDPD simulations and becomes even apparent during the calculation of the system properties after the system has reached equilibrium, but rather to assess the quality performance of the MDPD-MARTINI model in the simulation of DOPC and POPC bilayers with an additional C3 bead-type and investigate key properties of these bilayers. Typically, the simulated systems consist of 512 lipids, which are placed in a simulation box of dimensions $L_{x}=L_{y}=130$ and $L_{z}=100$, maintaining a 16:1 proportion of water beads.
@@
II.2 Many-body dissipative particle dynamics
The MDPD method has been applied for fluids with different properties [37, 28, 29, 30, 31, 32, 38, 39], as well as a range of soft matter multi-phase and multi-component systems, such as systems with surfactant molecules [40, 41]. The most common use of MDPD simulation is usually carried out by solving the Langevin equation of motion for each particle (Equation 1) as is also done in the case of MD simulations [42]. However, in the case of MDPD, forces are directly defined, instead of being provided by a potential form. The positions and velocities of each particle evolve according to Equation 1
@@ II.2 Many-body dissipative particle dynamics
which considers the interactions of a particular particle $i$ with all surrounding particles within a cutoff distance, $\bm{F}^{C}$, and the presence of dissipative, $\bm{F}^{D}$, and random forces, $\bm{F}^{R}$, to act as a thermostat. The mass $m$ is the same for all particles and set to unity.
The total conservative force on particle $i$, namely, $\bm{F}^{C}$, which is the main difference between dissipative particle dynamics (DPD) and MDPD [43, 44, 45, 46, 47] is mathematically expressed by the following relation
@@
The equations of motion (Equation 1) are integrated by a modified velocity-Verlet algorithm in LAMMPS [34, 35] with a time step $\Delta t=0.01$. The non-bonded interactions between beads are tuned by varying $A_{ij}$, while repulsive interactions will remain constant as $B=35$ [48]. In the MDPD model, harmonic bond and angle interactions are employed, which are mathematically expressed as
$E_{bond}=k\left(r_{ij}-r_{0}\right)^{2},$ (9)
@@
The MDPD model for the DPPC, POPC, and DOPC is illustrated in Figure 1. Moreover, the interactions between the different beads are presented in Table 2 [26]. Also, an additional bead-type C3 with respect to the DPPC lipid model has been considered as in the MD-MARTINI model in the case of the DOPC and POPC lipids. Finally, forming the bilayers by using the MDPD-MARTINI model follows the same simulation protocol as in the MD-MARTINI simulations with temperature increase from 300 K to 450 K. Typical snapshots for each bilayer case obtained from MDPD simulations are presented in Figure 2 alongside the corresponding MD-MARTINI snapshots.
@@
III Results and discussion
Figure 3: Comparison of bilayer thickness in MD (a, b) and MDPD (c, d) models. Results for POPC (a, c) and DOPC (b, d) bilayers constisting of 512 lipid molecules are shown. The values of bilayer thickness are reported in Table 4.
Lipid MD MDPD
DPPC 60.8 55.6
POPC 51.1 49.1
DOPC 49.8 52.6
Table 3: Area per lipid (APL) for each bilayer system obtained through MD and MDPD simulations in Å2.
Lipid MD MDPD
DPPC 42.5 42.8
POPC 40.3 41.7
DOPC 39.1 41.7
Table 4: Values of bilayer thickness for each lipid in Å.
The MD and MDPD MARTINI systems typically consisted of 512 lipids, while in certain cases, for example, for determining the spectrum intensities of the undulation modes on the bilayer, larger systems have been used (8192 lipds in total). Firstly, the area per lipid (APL) was examined. This can be determined by the area a lipid covers in such bilayers, in other words, this is the area of the bilayer divided by the average number of lipids in a single monolayer. In our study, the area of a plane aligned tangentially to the bilayer surface in the simulation box was divided by half the total number of lipid molecules in order to calculate the APL. This property depends on various factors, such as the lipid type and temperature. In this case, the influence of the lipid type was examined and a comparison between the MD and MDPD models was performed. The exact values of APL for DPPC, POPC, and DOPC bilayers are reported in Table 3. In particular, we can observed that APL values are in good agreement in both models, with the APL decreasing in the case of POPC and DOPC lipid bilayers. This is expected partly due to the different angle involving the C3-type beads (Table 1).
Another fundamental property that has been estimated is the thickness bilayer for the different lipids for both the MDPD- and MD-MARTINI models. This parameter is essential for describing structural aspects of the bilayers and was determined by idenftifying the peaks in the density distribution profile of the choline head-group bead (Q0) as shown in Figure 3. It is found that both the MD and MDPD models are in agreement with each other and moreover with experimental data [49] for all lipid bilayer cases. The exact values are reported in Table 4. In particular, it has been found that the thickness exhibits minor differences in the different systems. Only a slight decrease in the thickness is observed in the case of POPC and DOPC bilayers, while the structural characteristics expressed through the density profiles of each bilayer for the various bead types are similar for both the MD and the MDPD models, as shown in Figure 3.

### v2

@@ Abstract
(MDPD) offers a significant speed-up in the simulation of various systems, including soft matter, in comparison with molecular dynamics (MD) simulations based on Lennard-Jones interactions, which is crucial for describing phenomena characterized by large time and length scales. Moreover, it has recently been shown that the MARTINI force-field coarse-graining approach is applicable in MDPD, thus rendering feasible the simulation of complex molecules as in MD-MARTINI for ever larger systems and longer physical times. Here, simulations of various lipid membranes were performed by using the MDPD-MARTINI coarse-grained (CG) force-field (FF), relevant properties were calculated, and comparison with standard MD-MARTINI CG simulations and experimental data was made. Thus, insights into structural properties of these bilayer systems and further evidence regarding the transferability of the MDPD-MARTINI CG models are provided. In this regard, this is a natural next step in the development of the general-purpose MDPD-MARTINI-FF, which generally provides significant speed-ups in both computational and physical simulated times, in comparison with standard CG MD simulations.
Keywords:
@@
Many-Body Dissipative Particle Dynamics MDPD-MARTINI-FF Lipid Bilayers Water
journal: Eur. Phys. J. Plus
1 Introduction
@@
Molecular dynamics (MD) Rapaport_2004 is a common method used for the simulation of a whole range of materials in various areas, such as soft matter, fluid dynamics, and solid-state physics. A key element for reliable MD simulations is the ability of describing the interactions between particles (force field) in such a way that various experimental properties of interest of a system are well-reproduced by the simulation. To this end, the non-bonded interactions between particles play a key role in reproducing these properties and much of the research for creating the various force-fields (FFs) is mainly focused on adequately describing those non-bonded interactions for a given cutoff at a specific temperature (the development of FFs valid over a wide range of temperatures is an active field of research), with further refinements coming from the topology FF parameters, which often can also be crucial. When interactions are considered between all atoms of the system, that is in the case of all-atom FFs, simulations are generally computationally expensive, albeit often necessary, for example, when a finer time- and length-scale resolution is required, what has led to the development of coarse-grained (CG) FFs Noid2013 ; Kmiecik2016 ; Jin2022 ; Noid2023 . In this case, interactions between groups of atoms are modelled instead of individual atoms, thus the number of required force (interactions) calculations during the integration of the equations of motion in MD simulations is significantly reduced, which, in turn, results in faster simulations and significant savings in energy resources consumed to run the simulations and cool down the hardware. Moreover, CG models offer the possibility of using a larger time step in the simulation and faster dynamics, which provides an additional speed-up of the simulations.
@@
MARTINI is one of the most popular CG FFs for MD simulation, since it is a general-purpose FF that can be applied in the simulation of a wide spectrum of components Marrink2004 ; Marrink2007 ; Marrink2013 ; Alessandri2021 , such as proteins Monticelli2008 ; Periole2009 ; deJong2013 ; Poma2017 , polymers Lee2009 , carbohydrates Lopez2009 , glycolipids Lopez2013 , glycans Chakraborty2021 , DNA Uusitalo2015 , RNA Uusitalo2017 , water Yesylevskyy2010 and various solvents Vainikka2021 . In addition, various extensions include simulations for specific pH Grunewald2020 and more recently chemical reactions (reactive MARTINI) Sami2023 . While the MARTINI 3.0 CG FF Souza2021 has recently been released, the development of any FF is a continuous effort that aims at improving the FF parameters through refinement based on a wider range of methodologies and validation with experimental data on an ever growing number of possible systems. In the case of CG FFs, validation can often take place by comparison to results obtained by well-established all-atom FFs or CG models, which have already been parametrized according to experimental data.
@@
Regarding the MD-MARTINI FF, its popularity is mainly due to the fact that it allows for the simulation of a wide spectrum of different systems by only using a finite number of CG bead types and a top-down approach for non-bonded interactions. Here, the concept of a ’LEGO’ approach of combining these beads to build more complex molecules while accounting for chemical specificity is used. In addition, MARTINI is based on the Lennard-Jones potential with the smallest possible cutoff including attractions, which, in principle, renders this model computationally efficient, while various features implemented in GROMACS software (e.g. virtual atoms) Berendsen1995 , namely the native platform for MARTINI, allow for further speed-up of the simulations. Currently, the MD MARTINI method enjoys a large community, which applies and further tests the models in a range of diverse systems, thus practically participating in the validation effort of this general-purpose FF.
@@
Despite the wide use of the MD-MARTINI model, MD remains a computationally expensive method even for CG models. However, it has recently been shown that the MARTINI approach can be applied in the case of many-body dissipative particle dynamics (MDPD) Carnevale2024_MDPD_MARTINI , a method that offers a computational speed-up of about an order of magnitude in comparison with the MD method due to the short-range interactions, as well as an additional speed-up from the use of soft interactions, while at the same time it accounts for both repulsion and attraction between particles Pagonabarraga2001 ; warren2003 ; zhao2017 ; zhao2021-review ; zhao2021 ; Han2021 . In particular, it has been shown that the MDPD simulations are 4–7 times faster than the MD simulations for systems of DPPC lipids and water Carnevale2024_MDPD_MARTINI . Other comparisons in the literature have shown that MDPD models are four times faster than MD MARTINI models Zhu2021 , while in simpler systems (e.g. only water) MDPD models perform even faster Ghoufi2011 . Moreover, there is no limitation in combining MDPD and MD models in the same simulation. However, such a scenario is in general undesirable, since it will naturally lead to a higher computational cost for the simulation. Here, we build on previous work that dealt with the study of DPPC bilayers and further apply the MDPD-MARTINI-FF in the investigation of POPC and DOPC lipid bilayers and compare the obtained results with those of standard MD-MARTINI simulations. Such a comparison between the MD- and MDPD-MARTINI FFs is realized on the basis of a number of related properties for bilayer systems. Moreover, a new interaction level is provided to account for the C3 MARTINI bead-type in DOPC and POPC lipids. The MDPD-MARTINI models provide equivalent results to the MD MARTINI models, while they offer at the same time a significant speed-up in the simulations, namely several minutes of MDPD simulations instead of several hours of MD simulations, while interactions are transferable for these systems as in the case of the MD-MARTINI-FF.
@@
2 Simulation models and methods
Figure 1: MD and MDPD models of DPPC, DOPC, and POPC lipids. The type of beads are indicated by the different color. In particular, $Q_{0}$ represents the choline headgroup, $Q_{a}$ the phosphate headgroup, two $N_{a}$ beads correspond to the glycerol ester moiety, while the two alkane chains consist of $C_{1}$ and $C_{3}$ (for POPC and DOPC) beads. The POPC has a double bond at carbon atom positions 9 and 10, while the DOPC has a double bond at each chain, as illustrated. The interactions between the beads are given in Table 2.
2.1 Molecular Dynamics
The MD model follows that of previous work for DPPC bilayers Carnevale2024_MDPD_MARTINI , which has been a benchmark system in the case of the MD-MARTINI-FF lipids_tutorial and therefore further details will be here omitted. Initial files for LAMMPS software Plimpton1995 ; Thompson2022 for the DPPC, DOPC, and POPC bilayers based on the MD-MARTINI model can be found as moltemplate examples lipids_tutorial_moltemplate . The DPPC model consists of different bead types, that is $Q_{0}$ for the choline headgroup, $Q_{a}$ for the phosphate headgroup, two $N_{a}$ beads for the glycerol ester moiety, and two chains, each comprised of four $C_{1}$ beads, to represent the alkane tails. For DOPC and POPC lipids, the model differs from that of DPPC lipids. This distinction arises due to the presence of double bonds in the alkane chains, with one double bond in POPC and two in DOPC. These double bonds require a different angle in the chain structure and an additional bead type, C3, to accurately represent their molecular geometry. The lipid models are illustrated in Figure 1, with the names of the bead types being the same for both the MD and MDPD MARTINI models following the MARTINI FF naming convention. Hence, this implies that both models use the same mapping of CG beads to chemical groups of atoms. Finally, water is represented by the P4 bead type.
@@ 2.1 Molecular Dynamics
Typical simulations for forming the lipid bilayers follow the same protocol as in previous work Carnevale2024_MDPD_MARTINI . However, here, the aim is not to show the computational speed-up of the method during the bilayer formation, since this is well-established for MDPD simulations and becomes even apparent during the calculation of the system properties after the system has reached equilibrium Carnevale2024_MDPD_MARTINI , but rather to assess the quality performance of the MDPD-MARTINI model in the simulation of DOPC and POPC bilayers with an additional C3 bead-type and investigate key properties of these bilayers. Typically, the simulated systems consist of 512 lipids, which are placed in a simulation box of dimensions $L_{x}=L_{y}=130$ and $L_{z}=100$, maintaining a 16:1 proportion of water beads.
@@
2.2 Many-body dissipative particle dynamics
The MDPD method has been applied for fluids with different properties Espanol1995 ; warren2003 ; zhao2017 ; zhao2021-review ; zhao2021 ; Han2021 ; Vanya2018 ; Carnevale2023 ; Ng2025 , as well as a range of soft matter multi-phase and multi-component systems, such as systems with surfactant molecules Hendrikse2023 ; Carnevale2024 . The most common use of MDPD simulation is usually carried out by solving the Langevin equation of motion for each particle (Equation 1) as is also done in the case of MD simulations Theodorakis2011 . However, in the case of MDPD, forces are directly defined, instead of being provided by a potential form. The positions and velocities of each particle evolve according to Equation 1
@@ 2.2 Many-body dissipative particle dynamics
which considers the interactions of a particular particle $i$ with all surrounding particles within a cutoff distance, $\bm{F}^{C}$, and the presence of dissipative, $\bm{F}^{D}$, and random forces, $\bm{F}^{R}$, to act as a thermostat. The mass, $m$, is the same for all particles and set to unity.
The total conservative force on particle $i$, namely, $\bm{F}^{C}$, which is the main difference between dissipative particle dynamics (DPD) and MDPD Espanol2017 ; Yoshimoto2013 ; Li2014 ; Lavagnini2021 ; Trofimov2005 is mathematically expressed by the following relation
@@
The equations of motion (Equation 1) are integrated by a modified velocity-Verlet algorithm in LAMMPS Plimpton1995 ; Thompson2022 with a time step $\Delta t=0.01$. The non-bonded interactions between beads are tuned by varying $A_{ij}$, while repulsive interactions will remain constant as $B=35$ warren2013 . In the MDPD model, harmonic bond and angle interactions are employed, which are mathematically expressed as
$E_{bond}=k\left(r_{ij}-r_{0}\right)^{2}$ (9)
@@
The MDPD model for the DPPC, POPC, and DOPC is illustrated in Figure 1. Moreover, the interactions between the different beads are presented in Table 2 Carnevale2024_MDPD_MARTINI . Also, an additional bead-type C3 with respect to the DPPC lipid model has been considered as in the MD-MARTINI model in the case of the DOPC and POPC lipids. Finally, forming the bilayers by using the MDPD-MARTINI model follows the same simulation protocol as in the MD-MARTINI simulations with temperature increase from 300 K to 450 K. Typical snapshots for each bilayer case obtained from MDPD simulations are presented in Figure 2 alongside the corresponding MD-MARTINI snapshots.
@@
3 Results and Discussion
Figure 3: Comparison of bilayer thickness in MD (a, b) and MDPD (c, d) models. Results for POPC (a, c) and DOPC (b, d) bilayers consisting of 512 lipid molecules are shown. The values of bilayer thickness are reported in Table 3.
Table 3: MDPD and MD measurements of area per lipid and bilayer thickness for the different lipid molecules considered. Both methods produce membranes of comparable thickness, while the current MDPD force-field underestimates the APL.
MDPD MD
DPPC APL $46.2$2 $51.6$2
Thickness $44.2$ $44.0$
POPC APL $54.3$2 $61.6$2
Thickness $40.1$ $39.5$
DOPC APL $59.1$2 $64.7$2
Thickness $37.8$ $38.4$
The MD and MDPD MARTINI systems typically consisted of 512 lipids, while in certain cases, for example, for determining the spectrum intensities of the undulation modes on the bilayer, larger systems have been used (8192 lipids in total). Moreover, different sizes were tested in our previous work on DPPC lipids and all cases were able to form a stable bilayer Carnevale2024_MDPD_MARTINI . Firstly, the area per lipid (APL) was examined. This can be determined by the area a lipid covers in such bilayers, in other words, this is the area of the bilayer divided by the average number of lipids in a single monolayer. In our study, the area of a plane aligned tangentially to the bilayer surface in the simulation box was divided by half the total number of lipid molecules in order to calculate the APL. This property depends on various factors, such as the lipid type and temperature. In this case, the influence of the lipid type was examined and a comparison between the MD and MDPD models was performed. The exact values of APL for the lipid bilayers considered are reported in Table 3. In particular, we can observe that MDPD exhibits lower APL values than those from MD simulations. Moreover, both simulation methods present an increase in APL when going from DPPC–POPC–DOPC. This is expected partly due to the different angle involving the C3-type beads (Table 1).
Another fundamental property that has been estimated is the thickness bilayer for the different lipids for both the MDPD- and MD-MARTINI models. This parameter is essential for describing structural aspects of the bilayers and was determined by identifying the peaks in the density distribution profile of the choline head-group bead (Q0) as shown in Figure 3. It is found that both the MD and MDPD models are in agreement with each other and moreover with experimental data drabik2020 for all lipid bilayer cases. The exact values are reported in Table 3. In particular, it has been found that the thickness decreases as the APL increases for the different systems. The structural characteristics expressed through the density profiles of each bilayer for the various bead types are similar for both the MD and the MDPD models, as shown in Figure 3.